A banks loan officer rates applicants for credit. The ratings are normally distributed with a mean of 200 and a standard deviation of 50. if an applicant is randomly selected, find the probability of a rating that is between 200 and 275. Round to four decimal places.
Answer #1

Solution :

Given that ,

mean =   = 200

standard deviation = = 50   

P(200< x <275 ) = P[(200-200) / 50< (x - ) / < (275-200) / 50)]

= P(0 < Z < 1.5)

= P(Z <1.5 ) - P(Z < 0)

Using z table   

= 0.9332-0.5

probability= 0.4332